I have a question about how to set-up a rational exponent equation. the equation is L=2.43*W^0.3326. The W is 17Lbs. I am trying to figure out how to do the problem.
Answer by checkley77(12844)

W=17
L=2.43*W^0.3326
L=2.43*17^.3326
L=2.43*2.566
L=6.235 ANS.